OpenStudy (anonymous):

PLEASE HELP!!! Can someone help me with this problem? I need an exact answer for it! And if you could show me how to do it as well, that would be awesome! Thanks A BUNCH! :) A 2.9 kg mass starts from rest and slides down an inclined plane 8.3×10−1 m long in 0.56 s. What net force is acting on the mass along the incline? Answer in units of N

OpenStudy (anonymous):

how long is the plane? whats the angle of the incline?

OpenStudy (anonymous):

wait i see :)

OpenStudy (anonymous):

frictionless plane?

OpenStudy (anonymous):

i honestly have no idea! im assuming?

OpenStudy (anonymous):

F = ma so you need to find the acceleration it takes to slide. d = (1/2)a(t^2)

OpenStudy (anonymous):

if you solve the second equation for acceleration, then just plug that straight into the first equation

OpenStudy (anonymous):

wait im still a little iffy can you set the equations up for me?

OpenStudy (anonymous):

\[F = m(\frac{d}{\frac{1}{2}t^2}) = m\frac{2d}{t^2}\]

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

so the annswer would be 5.29?

OpenStudy (anonymous):

maybe if i did the equations right, and you typed it into the calculator right

OpenStudy (anonymous):

did anyone else get 5.29?

OpenStudy (ghazi):

 \[mg \sin \theta = m*a\] and \[s=0.5 *a*t^2\]now\[0.83= 0.5*a*(0.56)^2\] find a and get F